# Doppelgangers++: Improved Visual Disambiguation with Geometric 3D Features

### v1

@@ 2 Related Work
Disambiguation in SfM and image matching. Disambiguating similar structures and repeated patterns in SfM is a long-standing challenge. Prior work has mainly relied on heuristics-based analysis to detect conflicting relations and ambiguities in the structure of the underlying scene graph [23, 19, 20] or among image-level correspondence (or the lack thereof) [22, 7, 6]. While such methods have shown some success in detecting and resolving ambiguities in SfM pipelines, one fundamental limitation that they do not consider the rich information contained in the images themselves. In contrast, Doppelgangers [3] avoids handcrafted heuristics and takes a data-driven approach to the basic problem of identifying doppelganger image pairs. They train a CNN that classifies a pair of images as either positive or negative, with LoFTR [13]-extracted keypoints and match masks passed as auxiliary input. During inference, the binary classification results are used to pre-process the scene graph obtained from COLMAP to filter out spurious matches prior to running SfM. However, the method’s reliance on COLMAP feature extraction and matching and the need for auxiliary mask information from LoFTR introduces significant overhead and complexity to the overall pipeline, making it difficult to scale to larger scenes. In addition, we find that this prior method is brittle: it can fail to generalize to domains beyond landmarks (e.g., to office buildings), and it can require parameter tuning to get good results.
@@ 3.3 Evaluating Doppelganger correction in SfM
Specifically, for a given scene with known geolocation, we acquire from Mapillary a set of nearby geo-tagged images and register them to the reconstructed model with COLMAP [12]. We then apply RANSAC to robustly estimate a similarity transformation between the registered camera positions and their corresponding image geolocation metadata (converted to ECEF coordinates). We use the resulting RANSAC inlier ratio as an indicator of the model’s correctness. An example is shown in Fig. 4. We can observe that before correction, the registered cameras and their geolocations show significant misalignment, with the cameras collapsing to one side due to similar-looking surfaces. After correction, the registered cameras align closely to their true geolocations, indicating that doppelganger pairs are correctly separated in the model. If the model is split into multiple components, we calculate the weighted average of inliers ratios among the components:
@@ 4.1 Experimental setup
To comprehensively evaluate SfM reconstructions, we report the number of images in the SfM results, and the geo-alignment inlier ratio described in Sec. 3.3. Respectively, more images in the SfM model imply less false pair pruning, i.e., the model has better class separation; and a higher geo-alignment inlier ratio suggests that the reconstructed model is more likely to be accurate and complete. Note that the inlier ratio is also influenced by the accuracy of the geo-tagged images collected from Mapillary. Thus, if one test scene has more accurate geo-tagged images than another, it is likely to yield a higher inlier ratio. Therefore, rather than comparing inlier ratios across different scenes, this metric is more useful for evaluating different reconstruction methods on the same scene, as the same set of geo-tagged images is used.
@@ 4.4 Structure from Motion disambiguation
Table 2: Evaluation of SfM results. We compare reconstructions from COLMAP [12], DG-OG [3], and our method. $\tau$=0.8 is used unless otherwise stated. The ‘$+$’ symbol indicates split reconstruction components; e.g., DG-OG and our method split the Radcliffe Camera reconstruction into two components. Because VisymScenes scenes are large, we report statistics on the largest reconstruction component produced by COLMAP, and identify the corresponding (split) components in the results of DG-OG and ours.
@@ 6 Acknowledgment
We thank Joseph Tung and Brandon Li for their valuable discussion and help with the webviewer. This work was supported by the Intelligence Advanced Research Projects Activity (IARPA) via Department of Interior/Interior Business Center (DOI/IBC) contract number 140D0423C0035. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation thereon. Disclaimer: The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of IARPA, DOI/IBC, or the U.S. Government. This research was also supported in part by the National Science Foundation under award IIS-2212084.

### v2

@@ 2 Related Work
Disambiguation in SfM and image matching. Disambiguating similar structures and repeated patterns in SfM is a long-standing challenge. Prior work has mainly relied on heuristics-based analysis to detect conflicting relations and ambiguities in the structure of the underlying scene graph [35, 31, 32, 14, 18, 24, 27, 10, 19] or among image-level correspondence (or the lack thereof) [34, 13, 12, 20, 11, 4]. While such methods have shown some success in detecting and resolving ambiguities in SfM pipelines, one fundamental limitation that they do not consider the rich information contained in the images themselves. In contrast, Doppelgangers [3] avoids handcrafted heuristics and takes a data-driven approach to the basic problem of identifying doppelganger image pairs. They train a CNN that classifies a pair of images as either positive or negative, with LoFTR [25]-extracted keypoints and match masks passed as auxiliary input. During inference, the binary classification results are used to pre-process the scene graph obtained from COLMAP to filter out spurious matches prior to running SfM. However, the method’s reliance on COLMAP feature extraction and matching and the need for auxiliary mask information from LoFTR introduces significant overhead and complexity to the overall pipeline, making it difficult to scale to larger scenes. In addition, we find that this prior method is brittle: it can fail to generalize to domains beyond landmarks (e.g., to office buildings), and it can require parameter tuning to get good results.
@@ 3.3 Evaluating Doppelganger correction in SfM
Specifically, for a given scene with known geolocation, we acquire from Mapillary a set of nearby geo-tagged images and register them to the reconstructed model with COLMAP [23]. We then apply RANSAC to robustly estimate a similarity transformation between the registered camera positions and their corresponding image geolocation metadata (converted to ECEF coordinates). We use the resulting RANSAC inlier ratio as an indicator of the model’s correctness. As an example, in Fig. 4, we can observe that before correction, the registered cameras and their geolocations show significant misalignment, with the cameras collapsing to one side due to similar-looking surfaces. After correction, the registered cameras align closely to their true geolocations, indicating that doppelganger pairs are correctly separated in the model. If the model is split into multiple components, we calculate the weighted average of inliers ratios among the components:
@@ 4.1 Experimental setup
To comprehensively evaluate SfM reconstructions, we report the number of images in the SfM results, and the geo-alignment inlier ratio described in Sec. 3.3. Respectively, more images in the SfM model imply less false pair pruning, i.e., better class separation; and a higher geo-alignment inlier ratio suggests that the reconstructed model is more likely to be accurate and complete. Note that the inlier ratio is also influenced by the accuracy of the geo-tagged images collected from Mapillary. Thus, if one test scene has more accurate geo-tagged images than another, it is likely to yield a higher inlier ratio. Therefore, rather than comparing inlier ratios across different scenes, this metric is more useful for evaluating different reconstruction methods on the same scene, as the same set of geo-tagged images is used.
Baselines. Our primary baseline is [3], as their learning-based approach consistently outperforms heuristic-based methods like thresholding on the number of post-RANSAC inliers (e.g. SIFT [17]+RANSAC [9], D2-Net [8]+RANSAC, SuperPoint [6]+SuperGlue [22]) for pairwise visual disambiguation; and is also superior than previous approaches [12, 31, 5, 32] for SfM disambiguation.
@@ 4.4 Structure from Motion disambiguation
Table 2: Evaluation of SfM results. We compare reconstructions from COLMAP [23], DG-OG [3], and our method. $\tau$=0.8 is used unless otherwise stated. The ‘$+$’ indicates split reconstruction components; e.g., DG-OG and our method split the Radcliffe Camera reconstruction into two components. Because VisymScenes scenes are large, we report statistics on the largest reconstruction component produced by COLMAP, and identify the corresponding (split) components in the results of DG-OG and ours.
@@ 6 Acknowledgment
We thank Brandon Li and Joseph Tung for their valuable discussion and help with the webviewer. This work was supported by the Intelligence Advanced Research Projects Activity (IARPA) via Department of Interior/Interior Business Center (DOI/IBC) contract number 140D0423C0035. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation thereon. Disclaimer: The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of IARPA, DOI/IBC, or the U.S. Government. This research was also supported in part by the National Science Foundation under award IIS-2212084.